a body covers 5m in 1second and then covers10 m in 2 seconds find its average speed


1
Expert's answer
2021-05-04T12:08:05-0400

Average speed can be found as follows:


"v_{avg}=\dfrac{Total\ distance}{Time\ taken}."

Then, we get:


"v_{avg}=\dfrac{d_1+d_2}{t_1+t_2}=\dfrac{5\ m+10\ m}{1\ s+2\ s}=5\ \dfrac{m}{s}."